# Assurance of AI Systems From a Dependability Perspective

### v2

@@ 2.2 The Challenge of Assuring Perception
We have seen that an action subsystem can generally be assuredly guarded, but is dependent on a model of the world constructed by a perception subsystem that is harder to guard.
@@
A possible compromise is for the action guard to use a simpler, conservative model constructed by conventional and highly assured software. This is sometimes seen in self-driving cars where the guard is a simple system for automated collision detection and emergency braking, similar to the Automated Driver Assistance System (ADAS) that is provided for some human-driven cars. [^11] The US National Highway Traffic Safety Administration recently finalized a rule (FMVSS No. 127) that will require Automatic Emergency Braking (AEB), to be standard on all passenger cars and light trucks by September 2029. Tests by the American Automobile Association found that current AEBs (which are built to a lower, voluntary standard) are somewhat effective against rear-end collisions, but not at all effective against sideways collisions at intersections. However, this does not provide full assurance because forward collisions are not the only hazards (Mehmed and colleagues cite data from a NHTSA study of 5.9 million human-driver accidents classified in 37 categories [122]), nor are human accidents necessarily good models for failures of an autonomous system, and nor is emergency braking attractive as the sole method of hazard mitigation. And notice that to prevent excessive activation of the emergency system, the primary system of a self-driving car must take its behavior and capabilities into account. For example, the primary action subsystem may calculate (correctly) that a certain maneuver is optimal and safe, but that it is also likely to activate the more hair-trigger response of the emergency system, and so it must choose a different behavior. In general, systems employing defense in depth must be designed and developed as a whole to ensure that unintended differences among the layers do not cause unnecessary loss of availability, while safety arguments based on diversity are nevertheless preserved. These are challenging requirements.
@@
Furthermore, accidents and collisions are not the only hazards that should be considered: for example, the City of San Francisco has reported dozens of incidents where robotaxis interfered with emergency responders, and there must be many other circumstances where self-driving cars increase risk or inconvenience for others without themselves being involved in a collision.
@@
For verified dependability, assurable guards must detect all hazards and mitigate them safely, without excessive false alarms. It is possible that a more comprehensive suite of verifiable ADAS-like guard functions could do better, but they would have to steer a difficult path among incomplete coverage, false alarms, and unattractive, abrupt, responses. A contrary point of view is that although there may be many circumstances leading to accidents, the exact circumstances are irrelevant to “last second” detection as there are only a few possible emergency responses: essentially, braking or evasive action, and so an ADAS-like guard (or a suite of such guards) could be an acceptable means of assurance, provided we can develop assurance that it will always select and perform appropriate emergency responses (see Section 2.5).
@@
An argument against simple ADAS-like guards is that more sophisticated perception could detect hazardous situations earlier and provide less abrupt mitigation. Accordingly, it is worth considering guards that do use AI and ML and asking whether they can be assured for trustworthiness within an overall approach that remains close to the dependability end of the dependability/trustworthiness spectrum. For example, we could imagine an “ML-friendly” adjustment to traditional assurance where we do construct requirements and identify hazards, and then develop training data of “sufficient” size, coverage, and quality [^12] ML requires such a vast quantity of training data that the training sets are often labeled automatically, by another AI system, and therefore cannot be considered high quality. to encompass all of these and use it with a well-regarded ML toolset to generate the perception capabilities desired. In fact, this is exactly what is proposed by several groups engaged in research and development of autonomous systems [49, 19, 4], although they generally consider primary systems rather than guards, and only one focuses specifically on perception [154].
@@
Accurate perception for world models in general depends not only on robust interpretation of individual sensors but also on effective sensor fusion. This is illustrated by the fatal accident between an Uber self-driving car and a pedestrian walking a bicycle in Arizona on 18th March 2018 [125]. The Uber car used three sensor systems (cameras, radars, and lidar) and fused them using a priority scheme that delivered a “flickering” identification of the victim as the sensor systems’ own classifiers changed their identifications, and as fusion preferred first one sensor system, then another, as listed below [125, Table 1].
@@
We conclude that approaches based on trustworthy perception do not yet provide an adequate basis for assured world models, although they do represent good practice and can be welcomed and used on that basis.
@@ 2.3 Assurance through Diversity and Defense in Depth
The topic of assurance through diversity is large and somewhat contentious. There is little doubt that architectures employing diverse components are generally more reliable than single threads. In particular, there is evidence that “portfolio” or “ensemble” perception systems are more reliable than their individual constituents [87]. The difficulty is in demonstrating that diversity provides benefit in any particular case, and in estimating how much benefit it provides [111]. In particular, there is no feasible way to validate failure independence (it is a variant on the infeasibility of assurance by testing), [^13] However, it is feasible to validate modest degrees of independence, and indications of its absence. nor strong reasons for believing it. This is because some circumstances are just plain hard to interpret and it is possible that all components may then fail together: consider the scenario with the Cruise self-driving taxi described in Footnote 6—would any training set have included pedestrians being thrown under the wheels? Furthermore, these difficult cases do not “thin out” as more are considered: the distributions seem to have “fat tails” [98]. Thus, the “multiplicative” argument for assurance by diversity is indeed naive.
@@
However, although diversity alone cannot provide strong assurance, it can provide a useful step in a “ladder” of assurance. Nuclear power generation usually employs such a ladder of protection systems providing defense in depth: there is the operational control system, designed to manage the plant efficiently and safely, then a (safety) limitation system that can intervene to ensure the plant behavior stays operational but within some safe envelope, and finally a shutdown system that functions as an assured guard that guarantees to initiate a safe shutdown when safety parameters are violated. [^14] A complete shutdown can be a complex operation involving both automated initiation and intervention by expert operators, depending on the state of the plant and reasons underlying the need to shutdown (cf. the Fukushima incident). The operational and limitation systems are carefully engineered but do not guarantee the dependability goals established for nuclear power: that is accomplished by the assured shutdown system. But the operational and limitation systems are diverse in function and construction and although this does not provide strong assurance, the presence of the limitation system almost certainly reduces demands on the shutdown system. This is beneficial because an emergency shutdown is disruptive and expensive.
@@
The control, limitation, and shutdown systems in this architecture for nuclear power generation all use traditional software, but a similar approach could be used in systems with AI and ML. For example, in a self-driving car the full functionality could be provided by a primary system employing AI and ML that is supported by a diverse system, also employing AI and ML, that is focused on safety, with assurance provided by traditionally engineered ADAS-like emergency backup functions. The diverse primary and safety systems deliver acceptable behavior and reduce demands on the assured backup so that its interventions, though crude, are rare and tolerable. A safety system may introduce false alarms of its own, although these can be verifiably avoided in some circumstances [123]. Example AI safety functions include the “AI Safety Force Field” [126] that avoids creation of unsafe situations, and “REDriver” [164], which monitors proposed trajectories of self-driving cars against Chinese traffic laws.
@@
A criticism of these particular proposals for defense in depth is that the diverse safety-focused system is concerned solely with actions and relies on the same world model as the primary system. If the perception system fails to detect a pedestrian, for example, then it will be absent from the common model and the safety-focused system can deliver no protection: everything will depend on the emergency backup.
@@
One possible mitigation for this hazard is to provide the safety-focused system with separate sensors and perception; alternatively, it could use the same sensors as the primary system but with diverse perception software. Either arrangement would provide the primary and safety-focused systems with different models of the world (e.g., [70]) but it is debatable whether diverse models are manageable without false alarms and whether the ML-generated model for the safety system could be any more assurable than that for the primary. [^15] The model for the safety system could be simpler than that for the primary system. For example, the primary system in a self-driving car must not only detect objects, but their type (car, bicycle, pedestrian etc.) and fine details such as where drivers and pedestrians are looking, because the action system needs to project their likely motion and future position, whereas the safety system could just use objects and surround them with a conservative “safety box.” Intuitively, we might think that the simpler model of the safety system could be more trustworthy, but it would be difficult to provide credible assurance for this using current methods.
@@
It seems preferable to fuse the products of the diverse perception systems into a single model, but, as illustrated by the Uber crash described earlier, that can introduce flaws of its own. A more attractive approach is to construct a single world model using diverse perception systems in a principled way. Conventional perception systems work “bottom up”: one or more deep neural nets take sensor data (e.g., images from cameras or point clouds from lidars) and deliver interpretations (e.g., lists of detected objects) that are further processed and fused to produce the world model. One argument against this approach is that it works “backwards” from effect (image) to cause (objects), which is inherently difficult. Another is that it prioritizes fleeting sensor data above the world model, which is the repository of much accumulated information. An alternative approach, and the way human perception is believed to work [18], reasons “forwards” using the model to predict sensor data (or basic interpretations thereof) and then applies a form of Bayesian inference known as Variational Bayes [93] to optimize the model in a way that minimizes prediction error. Notice that prediction errors provide continuous feedback on accuracy of the world model.
@@
In humans, this mechanism for perception is known as Predictive Processing (PP) [176] and it is believed to be coupled with a dual-process architecture [60]. The lower-level process, known as “System 1” [89], performs rapid unconscious perception using PP so long as prediction errors are fairly small, indicating the world is evolving as expected. A large prediction error is called a “surprise” and the higher-level “System 2” process intervenes to resolve it using more deliberative cognition. For example, a self-driving car using a perception system of this kind may find that a vehicle some distance ahead that has been seen and correctly predicted for some time disappears, provoking a surprise; System 2 may hypothesize that the disappearance is due to it being occluded by another vehicle that has not been detected; System 2 can then add the hypothesized occluding vehicle to the model and it will be included in future predictions, thereby sensitizing basic sensor interpretation to look for it, and it can also initiate cautious defensive actions such as a lane change. [^16] This scenario is motivated by a Tesla crash in Taiwan on 1 June 2020: https://www.youtube.com/watch?v=ZmHBA_vV39w. System 2 in this dual-process architecture is a location for diverse refinement of the world model and also, optionally, the guarding of actions [85, 153]. Furthermore, minimization of prediction error provides a principled way to fuse diverse lower-level sensor and perception functions.
A PP perception pipeline can suffer systematic errors that cause it to misperceive some scenes. For a crude example, it could be blind to white trucks seen against clouds: sensors will report no such white trucks, leading to construction of a world model without white trucks; this will generate predictions lacking white trucks that are confirmed by the sensors. So all seems well—until we collide with a white truck. Suitable testing should minimize these circumstances [85], but diversity among sensors and their basic interpretation functions could also reduce them significantly.
@@
A dual-process architecture with System 1 using PP to integrate diverse perception systems and with System 2 providing diverse perception refinement and action guards is an attractive arrangement for autonomous systems. Although assurance for its AI and ML components will be lacking, its overall architecture has a rational structure that can justify modest assurance based on redundancy and diversity. Modest assurance is insufficient for system dependability, but that is not the claim it needs to support: instead, contributing to defense in depth, it supports the claim that it reduces demands on the traditionally engineered emergency backup to a tolerable level, [^17] For example, the primary system of a self-driving car might see an object ahead and classify it as a cardboard box and be prepared to drive over it; as it comes closer, the emergency backup will see it as an unclassified obstruction and slam the brakes on; but the safety system will also have seen it as an unclassified obstruction and could have changed lanes to avoid it “just in case.” On a larger scale, the safety system might veto a time-optimal route selection by the primary system because it considers it unduly hazardous (e.g., icy). and dependability is assured by the backup.
@@ 2.4 Summary of Architectural Choices
We divide the critical AI and ML components into a perception subsystem that builds a model of the local world and an action subsystem that uses the model to plan and execute safe and effective actions.
@@
    We begin at a point on the assurance spectrum where the AI and ML are considered “trustworthy.” This does not amount to credible assurance from the dependability perspective so both the world model the action subsystem are considered unassured. [^18] Those who wish to grant some confidence in the trustworthiness of well-examined AI and ML components may read our “unassured/unassurable” labeling as “somewhat assured.” “Assured” would then be read as “strongly assured,” with “weakly assured” as an intermediate level. Hence, we next focus on the dependability end of the spectrum.
@@
    As 3 but with diverse perception subsystems driving separate world models for the primary action subsystem and its guard. Diverse perception systems add little without cross-comparison (it could be the guard’s model that is flawed), so the guard is unassurable and the overall architecture reduces to 3 and is not assurable.
@@
    As 3 but with diverse perception systems contributing to a single world model. This is weakly assurable overall: the traditionally engineered guard provides assurance on actions but depends on a world model that is only weakly assured by a diversity argument on perception.

### v3

@@ 2.3 The Challenge of Assuring Perception
We have seen that an action subsystem can generally be assuredly guarded, but is dependent on a model of the local world constructed by a perception subsystem that is harder to guard.
@@
A possible compromise is for the action guard to use a simpler, conservative local model constructed by conventional and highly assured software. This is sometimes seen in self-driving cars where the guard is a simple system for automated collision detection and emergency braking, similar to the Automated Driver Assistance System (ADAS) that is provided for some human-driven cars (some of these themselves use ML-based perception—we consider this case later). [^12] The US National Highway Traffic Safety Administration recently finalized a rule (FMVSS No. 127) that will require Automatic Emergency Braking (AEB), to be standard on all passenger cars and light trucks by September 2029. Tests by the American Automobile Association found that current AEBs (which are built to a lower, voluntary standard) are somewhat effective against rear-end collisions, but not at all effective against sideways collisions at intersections. However, this does not provide full assurance because forward collisions are not the only hazards (Mehmed and colleagues cite data from a NHTSA study of 5.9 million human-driver accidents that used a classification with 37 categories [174]), nor are human accidents necessarily good models for failures of an autonomous system, and nor is emergency braking attractive as the sole method of hazard mitigation. And notice that to prevent excessive activation of the emergency system, the primary system of a self-driving car must take its behavior and capabilities into account. For example, the primary action subsystem may calculate (correctly) that a certain maneuver is optimal and safe, but that it is also likely to activate the more hair-trigger response of the emergency system, and so it must choose a different, suboptimal behavior. In general, systems employing defense in depth must be designed and developed as a whole to ensure that unintended differences among the layers do not cause unnecessary loss of availability, while at the same time striving to preserve safety arguments based on diversity. These are challenging requirements.
@@
Furthermore, accidents and collisions are not the only hazards that should be considered: for example, the City of San Francisco has reported dozens of incidents where robotaxis interfered with emergency responders, and there must be many other circumstances where self-driving cars increase risk or inconvenience for others without themselves being involved in a collision [141].
@@
For verified dependability, assurable guards must detect all hazards and mitigate them safely, without excessive false alarms. It is possible that a more comprehensive suite of verifiable ADAS-like guard functions could do better than a single safety system, but they would have to steer a difficult path among incomplete coverage, false alarms, and unattractive, abrupt, responses. A contrary point of view is that although there may be many circumstances leading to accidents, the exact circumstances are irrelevant to “last second” detection as there are only a few possible emergency responses: essentially, braking or evasive action, and so an ADAS-like guard (or a suite of such guards) could be an acceptable means of assurance, provided we can develop assurance that it will always select and perform appropriate emergency responses (see Section 2.6).
@@
An argument against simple ADAS-like guards is that more sophisticated perception could detect hazardous situations earlier and more completely, and provide less abrupt mitigation. Accordingly, it is worth considering guards that do use AI and ML and asking whether they can be assured for trustworthiness within an overall approach that remains close to the dependability end of the dependability/trustworthiness spectrum. For example, we could imagine an “ML-friendly” adjustment to traditional assurance where we do construct requirements and identify hazards, and then develop training data of “sufficient” size, coverage, and quality [^13] ML requires such a vast quantity of training data that the training sets are often labeled automatically, by another AI system, and therefore cannot be considered high quality. to encompass all of these and use it with a well-regarded ML toolset to generate the perception capabilities desired. In fact, this is exactly what is proposed by several groups engaged in research and development of autonomous systems [70, 26, 4], although they generally consider primary systems rather than guards, and only one focuses specifically on perception [215].
@@
Accurate perception for world models in general depends not only on robust interpretation of individual sensors but also on effective sensor fusion. This is illustrated by the fatal accident between an Uber self-driving car and a pedestrian walking a bicycle in Arizona on 18th March 2018 [179]. The Uber car used three sensor systems (cameras, radars, and lidar) and fused them using a priority scheme that delivered a “flickering” identification of the victim as the sensor systems’ own classifiers changed their identifications, and as fusion preferred first one sensor system, then another, as listed below [179, Table 1].
@@
We conclude that approaches based on trustworthy perception do not yet provide an adequate basis for assured local models, although they do represent good practice and can be welcomed and used on that basis.
@@ 2.4 Assurance through Diversity and Defense in Depth
The topic of assurance through diversity is large and somewhat contentious. There is little doubt that architectures employing diverse components are generally more reliable than single threads. In particular, there is evidence that “portfolio” or “ensemble” perception systems are more reliable than their individual constituents [122]. The difficulty is in demonstrating that diversity provides benefit in any particular case, and in estimating how much benefit it provides [161]. In particular, there is no feasible way to validate failure independence (it is a variant on the infeasibility of assurance by testing), [^14] However, it is feasible to validate modest confidence in independence, and to detect indications of its absence. nor strong reasons for believing it. This is because some circumstances are just plain hard to interpret and it is possible that all components may then fail together: consider the scenario with the Cruise self-driving taxi described in Footnote 7—would any test set have included pedestrians being thrown under the wheels? [^15] We would expect training to include other objects (e.g., traffic cones) being thrown under the wheels, and the hard common perception problem is then one of generalization. Furthermore, these difficult cases do not “thin out” as more are considered: the distributions seem to have “fat tails” [137] (see footnote 30 for an illustrative example). Thus, the “multiplicative” argument for assurance by diversity is indeed simplistic.
@@
However, although diversity alone cannot provide strong assurance, it can provide a useful step in a “ladder” of assurance. Nuclear power generation usually employs such a ladder of protection systems providing defense in depth: there is the operational control system, designed to manage the plant efficiently and safely, then a (safety) limitation system that can intervene to ensure the plant behavior stays within some safe envelope but remains operational, and finally a shutdown system that functions as an assured guard that guarantees to initiate a safe shutdown when safety parameters are violated. The operational and limitation systems are carefully engineered but do not guarantee the dependability goals established for nuclear power: that is accomplished by the assured shutdown system. But the operational and limitation systems are diverse in function and construction and although this does not provide a strong basis for assurance, the presence of the limitation system almost certainly reduces demands on the shutdown system. This is beneficial because an emergency shutdown is disruptive and expensive. [^16] A complete shutdown can be a complex operation involving both automated initiation and intervention by expert operators, depending on the state of the plant and reasons underlying the need to shutdown (cf. the Fukushima incident).
@@
The control, limitation, and shutdown systems in this architecture for nuclear power generation all use traditional software, but a similar approach could be used in systems with AI and ML. For example, in a self-driving car the full functionality could be provided by a primary system employing AI and ML that is supported by a diverse system, also employing AI and ML, that is focused on safety, with assurance provided by traditionally engineered ADAS-like emergency backup functions. The diverse primary and safety systems deliver acceptable behavior and reduce demands on the assured backup so that its interventions, though crude, are rare and tolerable. Example AI safety functions include “Responsibility-Sensitive Safety” (RSS) [218] and an “AI Safety Force Field” [181] that avoid creation of unsafe situations, and “REDriver” [228], which monitors proposed trajectories of self-driving cars against Chinese traffic laws.
@@
A criticism of these particular proposals for defense in depth is that the diverse safety-focused system is concerned solely with actions and relies on the same local model as the primary system. If the perception system fails to detect a pedestrian, for example, then it will be absent from the common model and the safety-focused system can deliver no protection: everything will depend on the emergency backup.
@@
One possible mitigation for this hazard is to provide the safety-focused system with separate sensors and perception so that it can build its own local model; alternatively, it could use the same sensors as the primary system but with diverse perception software where the local model for the safety system could be simpler than that for the primary system. For example, the primary system in a self-driving car must not only detect objects, but their type (car, bicycle, pedestrian etc.) and fine details such as where drivers and pedestrians are looking, because the action system needs to project their likely motion and future position, whereas the safety system could just perceive undifferentiated objects and surround them with a conservative “safety box.” Intuitively, we might think that the simpler model of the safety system could be more trustworthy, but it is challenging to provide credible assurance for this. Nonetheless, diverse sensors and/or diverse perception can provide the primary and safety-focused systems with different local models (e.g., [98]) and this should improve safety, but it is difficult to manage diverse models without false alarms (although these can be verifiably avoided in some circumstances [175]).
@@
Rather than manage two models, it seems preferable to fuse the products of the diverse perception systems into a single local model that combines precision and safety but, as illustrated by the Uber crash described earlier, fusing can introduce flaws of its own. A more attractive approach is to construct a single local model using diverse perception systems in a principled way. Conventional perception systems work “bottom up”: one or more deep neural nets take sensor data (e.g., images from cameras or point clouds from lidars) and deliver interpretations (e.g., lists of detected objects) that are further processed and fused to produce the local world model. One argument against this approach is that it works “backwards” from effect (image) to cause (objects), which is inherently difficult. Another is that it prioritizes fleeting sensor data above the local model, which is the repository of much accumulated information. An alternative approach, and the way human perception is believed to work [25], reasons “forwards” using the model to predict sensor data (or basic interpretations thereof) and then applies a form of Bayesian inference known as Variational Bayes [131] to optimize the model in a way that minimizes prediction error. Notice that prediction errors provide continuous feedback on accuracy of the local model. Furthermore, minimization of prediction error provides a principled way to fuse diverse lower-level sensor and perception functions.
@@
In humans, this mechanism for perception is known as Predictive Processing (PP) [243] and it is believed to be coupled with a dual-process architecture [83]. The lower-level process, known as “System 1” [124], performs rapid unconscious perception using PP so long as prediction errors are fairly small, indicating the world is evolving as expected. A large prediction error is called a “surprise” and the higher-level “System 2” process intervenes to resolve it using more deliberative cognition and a more comprehensive domain model. For example, a self-driving car using a perception system of this kind may find that a vehicle some distance ahead that has been seen and correctly predicted for some time disappears, provoking a surprise; System 2 may hypothesize that the disappearance is due to it being occluded by another vehicle that has not been detected (e.g., a white truck against clouds); System 2 can then add the hypothesized occluding vehicle to the local model and it will be included in future predictions, thereby sensitizing basic sensor interpretation to look for it, and the action system can also initiate cautious defensive actions such as a lane change. [^17] This scenario is motivated by a Tesla crash in Taiwan on 1 June 2020: https://www.youtube.com/watch?v=ZmHBA_vV39w. System 2 in this dual-process architecture is a location for diverse interpretation of the local model (and updates to resolve surprises) and also, optionally, the guarding of actions [118, 214].
@@
A dual-process architecture with System 1 using PP to integrate diverse perception systems and with System 2 providing diverse perception refinement and action guards is an attractive arrangement for autonomous systems. Although assurance for its AI and ML components will be weak, its overall architecture has a rational structure that can justify modest confidence in assurance based on redundancy and diversity. Modest confidence is insufficient for system dependability, but that is not the claim it needs to support: instead, contributing to defense in depth, it supports the claim that it reduces demands on the traditionally engineered emergency backup to a tolerable level, [^18] For example, the primary system of a self-driving car might see an object ahead and classify it as a cardboard box and be prepared to drive over it; as it comes closer, the emergency backup will see it as an unclassified obstruction and slam the brakes on; but the safety system will also have seen it as an unclassified obstruction and could have changed lanes to avoid it “just in case.” On a larger scale, the safety system might veto a time-optimal route selection by the primary system because it considers it unduly hazardous (e.g., potentially icy). and dependability is assured by the backup (the overall assurance case will integrate all these elements in its argument).
Even when a cyber-physical system does not employ a dual-process architecture, a second layer of runtime verification can have general benefit. The primary runtime verification layer checks for immediate safety hazards and intervenes to avoid or mitigate them, while the second layer checks for SPIs that act as “canaries in the mine” and warn of incipient or developing risks.
@@ 2.5 Summary of Architectural Choices
We divide the critical AI and ML components into a perception subsystem that builds a model of the local world and an action subsystem that uses the local model and its own domain model to plan and execute safe and effective actions.
@@
    We begin at a point on the assurance spectrum where the AI and ML are considered “trustworthy.” This does not amount to credible assurance from the dependability perspective so both the world model and the action subsystem are considered unassured. [^19] Those who wish to grant some confidence in the trustworthiness of well-examined AI and ML components may read our “unassured” category as “somewhat assured.” “Assured” could then be read as “strongly assured,” with “weakly assured” as an intermediate category, where the adjectives indicate confidence in the properties claimed for the components or system. Further variants would be those where the action subsystem uses no AI and is potentially assurable. Note that “assurable” concerns the ability to guarantee certain properties; some unassurable architectures may deliver generally good behavior, but we cannot guarantee it. Hence, we next focus on the dependability end of the spectrum.
@@
    As 3 but with diverse perception subsystems driving separate world models for the primary action subsystem and its guard. Diverse perception systems could provide benefit but without some cross-comparison the local models are not assured and neither is the guard and so the overall architecture reduces to 3 and is not assurable.
@@
    As 3 but with diverse perception systems contributing to a single world model that is considered weakly assured by a diversity argument. This architecture is weakly assurable overall: the traditionally engineered guard provides assurance on actions but depends on a world model that is only weakly assured.
@@
There are many variations on these architectures: for example, an assured backup guard could be added to Architecture 5 similar to Architecture 7, or the backup guard could be decomposed into perception and action subsystems giving rise to architectures with three “threads”: Doer, Checker, and Fallback [173]. Our goal is to encourage discussion of these topics, giving proper attention to the challenge of perception.